OpenStudy (anonymous):

A volume of 125mL of H 2 O is initially at room temperature (22.0 ∘ C ). A chilled steel rod at 2.00 ∘ C is placed in the water. If the final temperature of the system is 21.4 ∘ C , what is the mass of the steel bar? Use the following values: specific heat of water = 4.18 J/(g⋅ ∘ C) specific heat of steel = 0.452 J/(g⋅ ∘ C)

OpenStudy (anonymous):

The energy lost by the water = the energy gained by the steel rod, so: -q = q -mc(Tf-Ti) = mc(Tf-Ti) Just plug the values in and solve for mass of steel bar.

OpenStudy (anonymous):

Remember to change mL -> grams. Easy conversion.